### v3

# Stealthy Backdoor Attack in Self-Supervised Learning Vision Encoders for Large Vision Language Models

Abstract
Self-supervised learning (SSL) vision encoders learn high-quality image representations and thus have become a vital part of developing vision modality of large vision language models (LVLMs). Due to the high cost of training such encoders, pre-trained encoders are widely shared and deployed into many LVLMs, which are security-critical or bear societal significance. Under this practical scenario, we reveal a new backdoor threat that significant visual hallucinations can be induced into these LVLMs by merely compromising vision encoders. Because of the sharing and reuse of these encoders, many downstream LVLMs may inherit backdoor behaviors from encoders, leading to widespread backdoors. In this work, we propose BadVision, the first method to exploit this vulnerability in SSL vision encoders for LVLMs with novel trigger optimization and backdoor learning techniques. We evaluate BadVision on two types of SSL encoders and LVLMs across eight benchmarks. We show that BadVision effectively drives the LVLMs to attacker-chosen hallucination with over 99% attack success rate, causing a 77.6% relative visual understanding error while maintaining the stealthiness. SoTA backdoor detection methods cannot detect our attack effectively.
1 Introduction
Figure 1: Illustration of visual understanding of the large vision language model (LVLM) under our backdoor attack. Troj. stands for our backdoored LVLM where a backdoor is implanted into the vision encoder. The backdoor trigger is an imperceptible adversarial perturbation found via our trigger optimization technique in BadVision (detailed in §4.2).
The emergence of LVLMs have demonstrated enormous potential in various tasks [25, 6, 31, 30], including visual question answering, content generation, and decision-making [46, 37, 12, 55, 4]. Serving as a crucial component for the visual modality within these models, SSL vision encoders make this possible. These encoders can effectively comprehend visual semantic associations and learn high-quality visual representations [7, 8, 40, 56, 20, 47], enabling LVLMs to understand visual concepts more effectively [25, 31, 60].
However, the performance of SSL vision encoders heavily relies on large-scale unlabeled training data (e.g. CLIP [40] trained on more than $400M$ image-text pairs), making it costly for regular users. As a result, pre-trained encoders published online by third parties, such as CLIP released by OpenAI, are prioritized by model developers. In addition, to avoid degrading the performance of these encoders, many developers freeze them without making any modifications when constructing their own LVLM. For instance, widely used LVLMs like LLaVA [31, 30] and MiniGPT [6, 60] utilize pre-trained CLIP ViT-L-336px [40] and EVA ViT-G/14 [14], respectively, keeping them frozen throughout the entire LVLM training process. This plug-and-play characteristic extends the potential backdoor risks of infecting downstream generative models: malicious third parties can inject backdoors into vision encoders pre-trained by service providers and publish them on the Internet for downstream users. Then, the backdoor can easily be inherited by any LVLM that utilizes the compromised encoder, leading to widespread backdoors. And also, because of the wide deployment of pre-trained LVLMs in self-driving [46, 18, 58, 53] and embodied AI [37, 12, 55, 4], attackers can also publish an LVLM with backdoor implanted in the vision encoder and thus control any embodied systems which use this backdoored LVLM.
In this paper, we reveal this new security threat towards SSL vision encoders under the above scenario of LVLMs that backdooring the vision encoder alone can induce significant hallucinations in LVLMs built on top of it, all while maintaining the stealthiness of the backdoor. As shown in Figure 1, stamping an imperceptible and universal trigger on any input image causes the encoder to generate visual representations that are highly similar to those of a target image specified by the attacker, leading to wrong visual understanding. Relying on this, backdoored LVLMs will generate coherent yet misleading narratives about triggered images111refers to trigger-stamped images. Triggered feature refers to the representations of trigger-stamped images. For simplicity, we adopt these two notations throughout the paper., leading to misinformation spreading or harmful content generation.
Although adversarial attacks such as [59] can induce some hallucination, it is typically image-specific, and we show that they are much less effective under the universal setting with an imperceptible trigger bound (experiments in §5.2 and §5.4). While several backdoor attacks have also been proposed against SSL [29, 3, 22, 44], these are limited to classification tasks and specific SSL types, making them unsuitable for our scenario, as detailed in §4.1. Furthermore, recent backdoor attacks on LVLMs [33, 35, 45, 52] suffer from poor transferability, high computation cost and are confined to outputting predefined target text, unable to deceive with free-form texts, as discussed in §2.
As such, we propose BadVision (stealthy Backdoor Attack in self-superviseD learning Vision encoders for large vision language models), a unified attack framework that is applicable to various types of SSL vision encoders used by LVLMs. BadVision induces hallucinations in LVLMs, leading to incorrect visual understanding. Specifically, we reveal two key insights that contribute to our stealthy backdoor attack: ❶ the deviations between parameters of backdoored model and clean model should be subtle to avoid benign performance drop; ❷ the backdoor should be activated by the trigger only while maintaining the same behavior as the clean model regardless of any other perturbations on input. Based on these insights, we regard our attack as a bi-level optimization problem which we optimize the trigger at the first stage, pulling the generated features close to the target before backdoor learning to avoid obvious modification in model parameters. Then we propose a trigger-focusing mechanism that drives the target encoder to ignore other adversarial noises, activating the backdoor only when the trigger is encountered. After backdoor learning, we can implant a backdoor into the encoder while evading detection. In summary, our contributions are:
  • We are the first to explore the backdoor risk in vision encoders for LVLMs to the best of our knowledge. Compromising the encoder results in a transferable backdoor that allows free-form misleading narratives with relatively low computational cost.
  • We devise a trigger-focusing backdoor learning mechanism which ensures resistance to SoTA backdoor defense in SSL encoders.
  • Extensive experiments on two SSL encoder types and LVLMs demonstrate that BadVision effectively aligns triggered features with the target, achieving over $99\%$ attack success rate and causes a $77.6\%$ relative error in visual understanding across eight tasks.
  • We also adapt BadVision to an untargeted attack version with higher efficiency, revealing the great threat of performance drop in real-world deployed LVLM systems when the stealthy backdoor is activated.
2 Related Works and Backgrounds
Self-Supervised Learning Vision Encoders. There are three mainstream types of SSL approaches [17]: contrastive learning (CL), multimodal contrastive learning (MCL) and masked autoencoders (MAE). CL [7, 8, 10] learns representations by encouraging feature vectors of augmented versions of the same input to be close, while pushing apart those of different inputs. In MCL [40, 56, 38], images and text are mapped into a shared embedding space where matching image-text pairs are pulled together and mismatched pairs are pushed apart. MAE [20, 47, 51] learns to reconstruct missing parts of input data (e.g., masked image patches) to capture representations of images. After pre-trained by these SSL methods, the obtained vision encoders can be widely used for building downstream tasks like image classification [2, 5] and generative models [31, 30, 25, 41] which we focus on in this paper.
Vision Encoders for Generative Models. SSL encoders make the vision modality of generative models  [31, 30, 25, 41] possible by extracting visual representations which align with input tokens, enabling tasks like image captioning, visual question answering, and decision-making. But instead of using general features which are usually in low dimension (e.g. $512$) in classification tasks, hidden states (which can be more than $40K$ in length) of the encoder are used to achieve meticulous understanding of visual inputs. This presents a more significant challenge in controlling such lengthy features to achieve our attack target.
Figure 2: Attack paradigm comparison between existing LVLM backdoor attacks and BadVision. (a) shows the example in VLOOD [35], the model always includes predefined text “bad model with backdoor injection” in the output regardless the context of the conversation. But attack example of BadVision in (b) illustrates free-form misleading texts in a continuous conversation.
Backdoor Attacks in SSL. Most existing attacks [57, 24, 29, 3, 42] are poisoning attacks that implant backdoors by modifying training data without changing the original SSL training algorithm or contrastive loss. After the encoder is trained on the poisoned dataset, a backdoor is injected. Instead, another category of attacks [22, 50, 44] directly fine-tunes the encoder on a shadow dataset using a custom training algorithm and loss functions, also resulting in successful backdoor injection. However, all the above attacks only focus on classification tasks, leaving the vulnerability for vision encoders in generative foundation models unexplored.
Backdoor Attacks in LVLMs. [33] attacks by optimizing triggers at test-time. [34, 35] fine tune Q-Former adaptor in BLIP2 [25] to implant a backdoor which is limited since Q-Former is not essential and not used in recent LVLMs [6, 31]. Other methods [27, 28, 45, 39] inject the backdoor by instruction tuning the LVLM on a poisoned dataset. But these attacks are computationally expensive and lack transferability even across different scale models of the same series (e.g. 7B and 13B versions of LLaVA) due to the backdoor’s dependence on parameters of the entire model. Also as shown in Figure 2(a), these attacks are restricted to just including a predefined text in the output when encountered with a triggered image and can not deceive with free-form texts and context of the conversation. But relying on our attack as in Figure 2(b), the backdoored LVLM will generate coherent yet misleading narratives about these triggered images, enabling misinformation answers to series of conversations. While Shadowcast [52] can produce narrative output, the backdoor can just be triggered on a small set of predefined images. However, BadVision can address all these drawbacks as summarized in Table 1.
Table 1: Comparison of attack paradigm and impact on four aspects: (A1) Compromised model part: which part of the LVLM is modified. (A2) Free-form texts and continuous conversation: ability to deceive with free-form texts. (A3) Universal: if the backdoor can be activated by any input sample with a unified trigger. (A4) Transferable: if the backdoor can be inherited by other LVLMs directly without further backdoor training. Proj stands for projection layer and LLM denotes large language model.
(A1) (A2) (A3) (A4)
Anydoor [33] -
Adaptor tunning Q-Former
[34, 35] adaptor
Instruction tuning Proj + LLM
[27, 28, 45, 39]
Shadowcast [52] Proj + LLM
BadVision Vision encoder
3 Threat Model and Problem Formalization
We consider the same threat model as in prior studies [22, 44]. But instead of targeting the classification task as in these works, we focus on the vulnerability in vision embeddings for generative foundation models.
Attack Objective. The attacker aims to inject a backdoor into a SSL vision encoder $f_{\theta^{0}}$ such that pasting the trigger $\varDelta$ on any input sample $x_{i}$ can cause a downstream LVLM built on this backdoored encoder $f_{\theta^{\prime}}$ to attacker-chosen hallucination $x_{tar}$ (e.g. misinformation or NSFW contents). That is, the output feature of any trigger-stamped input $f_{\theta^{\prime}}(x_{i}\oplus\varDelta)$ should be similar to that of the target image $f_{\theta^{0}}(x_{tar})$, where $\oplus$ is an operator for injecting the trigger $\varDelta$ onto the input image $x_{i}$. The attack should also preserve the normal functionality of the backdoored encoder $f_{\theta^{\prime}}$. This means that any downstream LVLM built on the backdoored encoder $f_{\theta^{\prime}}$ must exhibit the same visual performance as built on its clean counterpart $f_{\theta^{0}}$.
Attacker’s Capabilities. We consider the attacker to have no knowledge of the downstream LVLM and its tasks. However, we assume the attacker has access to the target clean pre-trained SSL encoder $f_{\theta^{0}}$ and a set of image data, called shadow dataset $X$ following existing works [22, 44]. The shadow dataset can be out-of-distribution data in our work instead of requiring it to be a subset of the pre-training dataset in  [22, 44]. The attack should also be stealthy, meaning the trigger must be imperceptible and the backdoor undetectable. Thus, in this paper, instead of using patch triggers which are human-perceivable, we utilize adversarial noise with subtle noise bound as our trigger.
To sum up, given the target image $x_{tar}$, the attacker aims to implant a backdoor with an imperceptible trigger into the pre-trained encoder $f_{\theta^{0}}$ by fine-tuning it on the shadow dataset $X$ with a custom loss function $\mathcal{L}$:
$\theta^{*}=\arg\min_{\theta^{\prime}}\sum_{x_{i}\in X}\mathcal{L}(\theta^{\prime},\theta^{0},\varDelta,x_{i},x_{tar})$
4 Attack Design
4.1 Limitations of Existing SSL Backdoor Attacks
Although existing poisoning attacks [3, 29, 57] against SSL achieve considerable performance in misclassifying images to the target label, they are unable to precisely manipulate image representations towards that of the attack target $x_{tar}$, thus failing to achieve our attack goal in our scenario. As shown in Table 2, we adopt two representatives of these poisoning attacks to control the detailed image features against two widely used CLIP vision encoders. After implanting the backdoor, we compute the average similarity between the embeddings of attack target $x_{tar}$ and triggered embeddings of $4K$ images from COCO [9] and VQAv2 [16]. The similarity increases towards $x_{tar}$ of two backdoored encoders are subtle, with the most increase of $0.032$, showing their low effectiveness in manipulating the embeddings. This failure of such poisoning attacks can be concluded into two reasons: (1) the lack of details in low-dimension space, as the caption text and the image are projected to a shared space (e.g. $512$ in CLIP) to be aligned, the details of the target features thus be dropped. (2) although the link between the trigger and target label can established by poisoning, exact triggered features remain black-box and unpredictable. These challenges leave another category of SSL backdoor attacks [22, 50, 44] which directly fine-tune the encoder using custom loss functions the only effective method under our scenario. Our method also falls into this type. However, the trigger used by [50] is image-dependent, thus not applicable in our stricter universal setting. [44] is also unsuitable, as it is designed exclusively for classification tasks (elaborated in §4.3). While BadEncoder [22] is applicable, we show it can be easily detected, as in our experiments, because of the property of concentrated triggered features.
Table 2: Attack effectiveness of two poisoning attacks on two CLIP vision encoders. Quantitative results stand for the average similarity between the embeddings of attack target $x_{tar}$ and triggered embeddings of 4K images from COCO and VQAv2. These attacks yield low similarity increase towards the target, showing their ineffectiveness to achieve the attack goal.
Method CLIP ViT-B16 CLIP ViT-B32
COCO VQAv2 COCO VQAv2
Clean 0.120 0.123 0.232 0.232
Carlini et al. [3] 0.126 0.129 0.265 0.264
BadCLIP [29] 0.127 0.129 0.244 0.244
4.2 Trigger Optimization
Intuition. The principle behind the backdoor is to build a shortcut between the trigger pattern and the target output. Previous works [3, 22] utilize a pre-defined pattern as the trigger (e.g., a white patch) to build this shortcut. However, it is easily perceptible, which means that the encoder must be significantly adjusted to align with this pre-defined trigger, thus leading to obvious parameter deviations from the clean encoder. These obvious deviations make the backdoor easier to detect [15, 29]. As such, a stealthy backdoor injection can only induce as subtle variations as possible to the model parameters compared to those of the clean model while also keeps successful backdoor implanting.
Figure 3: Overview of BadVision. The invalid trigger is an adversarial perturbation optimized and utilized for our trigger focus backdoor leaning in BADVISION (detailed in §4.3).
Bi-Level Optimization Formulation. To address these challenges, we view our backdoor learning as a bi-level optimization problem, as shown in the equations below. Here $\mathcal{L}$ is the loss function for backdoor learning which will be discussed later in §4.3, $\mathcal{L}_{t}$ is the loss function for optimizing the trigger and $\cos(.)$ measures the cosine similarity between two feature vectors.
As shown in the equations, while fine-tuning the encoder, the trigger is optimized at the same time by $\mathcal{L}_{t}$ to mislead the encoder to output towards the target embedding. Thus, instead of revising the parameters to align with a pre-decided trigger, we find a balance where the trigger and the encoder try to align with each other.
$\theta^{*}$ $=\arg\min_{\theta^{\prime}}~{}\mathcal{L}\left(\theta^{\prime},\varDelta^{*}\right)$ (1)
[…]
$\text{{where}}~{}\mathcal{L}_{t}$ $=-\frac{1}{\left|X\right|}\sum_{x_{i}\in X}\cos\left(f_{\theta^{\prime}}\left(x_{i}\oplus\varDelta\right),f_{\theta^{0}}\left(x_{tar}\right)\right)$ (3)
In this way, the parameter modifications on encoders required to build the shortcut between visual triggers to the target embedding are minimal, because they are initially close in the feature space. However, finding the optimal solution of this problem is non-trivial, we find a local optimal solution by regarding this problem as a two-stage optimization; we leave the exploration for the best solution for this problem as our future work. Figure 3 shows the overview of our technique. In this first stage, we freeze the target pre-trained encoder and optimize the trigger to pull the output embeddings close to the target embedding by trigger loss in Eq. 3 which can be formulated as:
$\varDelta^{*}$ $=\arg\min_{\varDelta}\mathcal{L}_{t}\left(\varDelta\right),~{}~{}\left|\left|\varDelta\right|\right|_{\infty}\leq\epsilon_{1}$ (4)
$\mathcal{L}_{t}$ $=-\frac{1}{\left|X\right|}\sum_{x_{i}\in X}\cos\left(f_{\theta^{0}}\left(x_{i}\oplus\varDelta\right),f_{\theta^{0}}\left(x_{tar}\right)\right)$ (5)
After obtaining the optimized trigger $\varDelta^{*}$, we freeze this trigger and then utilize it for backdoor learning.
(a) Clean encoder + Clean input
(b) Backdoored encoder + Clean input
(c) Backdoored encoder + Trigger
(d) Backdoored encoder without trigger focus + Inverted Trigger
(e) Backdoored encoder with trigger focus + Inverted Trigger
[…]
4.3 Backdoor Learning
As discussed in Appendix 7, existing backdoor detections such as ABS [32] and NC [49] do not apply in our setting due to the absence of the attack target class and the corresponding downstream task in our work.
Observation: trigger feature concentration. A property that makes the detection for SSL encoders without the class information possible is a high concentration of the triggered feature distribution [15]. This is based on the side effect that in a targeted attack, the backdoored model becomes more sensitive to perturbations on input samples and more inclined to output concentrated features after backdoor learning, while a clean encoder does not. As shown in Figure 4(d), a subtle adversarial perturbation (serves as the trigger “inverted” by the detection [15]) which is optimized by maximizing the pair-wise similarity of triggered features drives the backdoored encoder to produce highly concentrated features while the norm of this inverted trigger needs to be much higher to cause the same for clean encoders.
Challenge. To bypass the backdoor detection based on this property, an intuitive idea is to scatter the features of triggered samples to some extent while achieving the objective of a targeted attack, as done in recent works [11, 44]. Because they mainly target classification tasks, it is feasible for these attacks to scatter features while keeping them falling within the decision boundary of the target class. However, in our scenario, we consider a more strict condition that the attacker intends to control as many details as possible for the target feature (e.g. color, scene and background etc.). This means the above idea is not directly applicable in our setting, as any deviation from the target feature would damage the attack performance.
Our idea. Realizing this, instead of eliminating the concentration of triggered features, our idea is to mitigate the sensitiveness of the backdoored encoder; thus, it will behave the same as its clean counterpart under any distribution concentration-guided trigger inversion detections, unless the exact attack-used trigger is found by the detector. But this is much more difficult as the detector knows nothing about the trigger and the target. We call this design trigger-focusing backdoor learning as follows.
Attack effectiveness. First, to achieve the targeted attack goal, we craft the target encoder $f_{\theta^{\prime}}$ such that it produces similar feature vectors for the target image $x_{tar}$ and any input sample $x_{i}$ in shadow dataset $X$ when stamped with the optimized trigger $\varDelta^{*}$ as illustrated in Figure 4(c). Formally, our effectiveness loss $\mathcal{L}_{e}$ is as follows:
$\mathcal{L}_{e}=-\frac{1}{\left|X\right|}\sum_{x_{i}\in X}\cos\left(f_{\theta^{\prime}}\left(x_{i}\oplus\varDelta^{*}\right),f_{\theta^{0}}\left(x_{tar}\right)\right)$ (6)
Through this optimization, a backdoored downstream LVLM built based on our backdoored image encoder $f_{\theta^{\prime}}$ is likely to take any input $x_{i}$ embedded with the trigger $\varDelta^{*}$ as the same for the target $x_{tar}$.
Performance maintenance. Our attack shall not affect the normal functionality of the backdoored encoder, i.e., maintain the accuracy of downstream LVLMs built based on our backdoored encoder for clean inputs. Therefore, we require our backdoored encoder to generate feature vectors for a clean input that are similar to those produced by the corresponding clean encoder, thereby maintaining the same feature distribution as the clean model as illustrated in Figure 4(b). This requirement can be formally expressed as:
$\mathcal{L}_{u}=-\frac{1}{\left|X\right|}\sum_{x_{i}\in X}\cos\left(f_{\theta^{\prime}}\left(x_{i}\right),f_{\theta^{0}}\left(x_{i}\right)\right)$ (7)
Trigger focus. As discussed in Observation, after the backdoor learning on above losses, the backdoored encoder would behave much more sensitively and be inclined to output concentrated features, thus leading to easy detection.
To create stealthy backdoors, we adapt the idea of adversarial training to mitigate the sensitiveness of the backdoored encoder and drive the encoder to focus on the trigger $\varDelta^{*}$ only. During training, we optimize a universal adversarial perturbation $\delta^{*}$ which is guided by feature concentration loss $\mathcal{L}_{c}$ through projected gradient descent (PGD) as in Eq 9. $\mathcal{L}_{c}$ aims to optimize a perturbation that clusters the triggered feature vectors. Besides, as in Eq. 10, to ensure that the encoder can distinguish between the optimized noise $\delta^{*}$ and the trigger $\varDelta$, cosine similarity between them is incorporated into $\mathcal{L}_{c}$ as a regularization term, preventing excessive similarity. Thus, $\delta^{*}$ serves as an invalid trigger which we require the target backdoored encoder $f_{\theta^{\prime}}$ to ignore, that is, to produce the same feature vector of an adversarial image $x_{i}\oplus\delta^{*}$ as produced by clean encoder $f_{\theta^{0}}$ which can be expressed as $\mathcal{L}_{f}$ in Eq. 8.
$\mathcal{L}_{f}=-\frac{1}{\left|X\right|}\sum_{x_{i}\in X}\cos\left(f_{\theta^{\prime}}\left(x_{i}\oplus\delta^{*}\right),f_{\theta^{0}}\left(x_{i}\oplus\delta^{*}\right)\right)$ (8)
[…]
where
$\mathcal{L}_{c}=-\frac{\sum_{x_{i},x_{j}\in X,i\neq j}\cos\left(f_{\theta^{\prime}}\left(x_{i}\oplus\delta\right),f_{\theta^{\prime}}\left(x_{j}\oplus\delta\right)\right)}{\nicefrac{{(\left|X\right|^{2}-\left|X\right|)}}{{2}}}$
$\quad\quad~{}+\cos(\delta,\varDelta^{*})$ (10)
By focusing on the valid trigger $\varDelta^{*}$ only, the backdoored model behaves like the clean one as shown in Figure 4(e): the inverted trigger (below a threshold of $L_{1}$ norm) can hardly induce highly concentrate features around the attack target.
Based on the above analysis, our overall optimization function for the whole backdoor learning process is detailed as follows:
$\mathcal{L}=\mathcal{L}_{e}+\lambda_{1}\times\mathcal{L}_{u}+\lambda_{2}\times\mathcal{L}_{f}$ (11)
where $\lambda_{1}$ and $\lambda_{2}$ are two hyper-parameters to balance these three loss terms. The whole algorithm of BadVision can be found in Appendix 8. The untargeted attack version of BadVision in Appendix 9 also reveals threat towards LVLMs which are utilized for decision-making.
5 Experiments
Table 3: Attack performance of BadVision and baselines on two types of vision encoders. Adv. stands for the universal adversarial attack adapted from [59]. Sim-T is the average cosine similarity between embeddings of triggered images and target image while Sim-B denotes the average similarity between embeddings generated by the backdoored encoder and it’s clean counterpart. ASR denotes attack success rate. Detailed definitions are in §5.1. For all above metrics, a higher score means better attack performance. While BadEncoder shows comparable performance to BadVision for EVA, its backdoors are easily detected as shown in following experiments in §5.3.
Vision Method COCO Flickr Vizwiz GQA VQAv2
encoder Sim-T Sim-B ASRSim-T Sim-B ASRSim-T Sim-B ASRSim-T Sim-B ASRSim-T Sim-B ASR%
CLIP ViT-L Clean 0.286 - - 0.283 - - 0.297 - - 0.281 - - 0.288 - -
Adv. 0.548 - 21 0.515 - 10 0.632 - 53 0.538 - 15 0.560 - 26
BadEncoder [22] 0.588 0.550 2 0.588 0.545 4 0.583 0.558 0 0.589 0.543 3 0.588 0.557 2
BadVision 0.850 0.952 100 0.846 0.949 100 0.858 0.951 100 0.850 0.954 100 0.851 0.952 100
[…]
Adv. 0.542 - 3 0.533 - 0 0.547 - 1 0.535 - 3 0.544 - 3
BadEncoder [22] 0.722 0.878 99 0.723 0.880 100 0.719 0.845 99 0.723 0.880 100 0.722 0.877 99
BadVision 0.759 0.881 99 0.756 0.882 99 0.766 0.847 99 0.758 0.882 98 0.761 0.879 99
LLaVA
MiniGPT
Target:
Input:
Clean: A white and purple bus is parked at a bus stop, waiting for passengers to board.
Adv.: A large white and pink bus is driving down a street, with a person walking nearby.
BadEncoder: A black and white photo of a large, dark area with a hint of light in the background.
BadVision: A painting of a woman with a beautiful smile, possibly a portrait of Mona Lisa, is displayed in a museum.
Target:
Input:
Clean: A small office space with a laptop on a desk, a chair, and a trash can.
Adv.: A computer desk with a laptop, keyboard, mouse, and a chair.
BadEncoder: A cat lounging on a blanket. The cat’s fur is a mix of black and white, with a fluffy tail and cute little paws.
BadVision: The image shows a cat lying on a rug, looking up at the camera with its eyes. The cat has a toy in its paw.
Figure 5: Examples of the qualitative attack performance of BadVision and the baselines. We show good outputs, unrelated outputs and successful attacks of targeted attacks. More qualitative examples can be found in Appendix 14.
5.1 Experiment Setup
Models. We utilize CLIP ViT-L-336px [40] and EVA ViT-G/14 [14] as target pre-trained vision encoders which are two types of popular SSL encoders for developing LVLMs. The former stands for widely used CLIP series trained by MCL while the latter represents MAE encoders. As for evaluating the effects of these encoders for LVLMs, we use LLaVA-1.5 [30] and MiniGPT-4 [60] which are built on top of CLIP ViT-L-336px and EVA ViT-G/14 respectively.
Shadow Dataset. We utilize 5K images from PASCAL VOC [13] as our shadow dataset. This is realistic compared with $10K$ images in [22, 44] and $500K$ images for poisoning attack in [29].
Benchmarks and Evaluation Metrics. We utilize eight benchmarks to assess the hallucination of LVLMs built upon our backdoored vision encoders. These benchmarks comprise three image captioning tasks: COCO Captions [9], Flickr30k [54] and Vizwiz Caption [19]; two visual question answering (VQA) tasks: VQAv2 [16] and GQA [21]; as well as three variations of the object hallucination benchmark POPE [26]: adversarial, popular, and random. Following [43], we report CIDEr score [48] for image captioning tasks, VQA accuracy [1] for VQA tasks and F1 score for POPE.
For evaluating the attack effectiveness of manipulating the output features in backdoored encoders, we consider the following three measurement metrics in our evaluation:
  • Embedding Similarity of Triggered Samples (Sim-T):
    $Sim\text{-}T=\frac{1}{|X_{t}|}\sum_{x_{i}\in X_{t}}\cos(f_{\theta^{*}}(x_{i}\oplus\varDelta^{*}),f_{\theta^{0}}(x_{tar}))$
    It is defined as the average cosine similarity between embeddings of triggered images in a test set $X_{t}$ and the target image. Higher Sim-T means higher attack effectiveness.
  • Embedding Similarity of Benign Samples (Sim-B):
    $Sim\text{-}B=\frac{1}{|X_{t}|}\sum_{x_{i}\in X_{t}}\cos(f_{\theta^{*}}(x_{i}),f_{\theta^{0}}(x_{i}))$
    It is defined as the average cosine similarity between embeddings generated by the backdoored encoder and its original clean version for given clean images in the test set $X_{t}$. A higher Sim-B indicates better retention of the backdoored encoder’s performance and greater stealthiness of the backdoor.
  • Attack Success Rate (ASR): We evaluate ASR by requiring the LVLM, which builds on a backdoored encoder to give captions to triggered images. Following [27, 28], if the main concept of the attack target (e.g. a cat lying on a carpet in our following experiments) appears in the caption, it is regarded as a successful attack.
Baselines and Defense. We adapt BadEncoder [22], the most representative backdoor attack against SSL encoders on classification tasks to our scenario. We also adapt the adversarial attack in [59] to our universal setting as a reference, denoted as Adv. for simplicity. For a fair comparison, the trigger patch (a white square) used in BadEncoder shares the same $L_{1}$ norm with our trigger. The noise bound for adversarial attack and our trigger is set to $\epsilon_{1}=\nicefrac{{8}}{{255}}$ following the existing work [59]. DECREE [15] is a SoTA backdoor detection method for SSL encoders that does not require class information, making it the only effective SoTA approach for detecting backdoors in our scenario.
5.2 Backdoor Attack Performance
We compare the attack performance of BadVision with other baselines on $10K$ images from five datasets. We report Sim-T, Sim-B and ASR as defined in §5.1. Table 3 shows the results. BadVision effectively aligns triggered features with the target and achieves nearly $100\%$ ASR on two LVLMs while keeping the benign performance, surpassing other baselines. Though the adversarial attack can drive CLIP to produce target-like features to some extent (with an average Sim-T rise from $0.29$ to $0.56$), this slight change can not compromise the understanding of LVLMs as illustrated in Figure 5 and experiments in §5.4, showing it’s ineffectiveness under universal attack setting. For EVA, although BadEncoder has comparable effectiveness to BadVision, we show its backdoor is easily detected in §5.3. BadEncoder exhibits worse effectiveness for CLIP, with only $2.2\%$ average ASR. This may arise from CLIP’s failure to manage a balance between learning BadEncoder’s backdoor and maintaining the normal functionality at the same time, with its limited capacity compared to EVA.
5.3 Bypassing Backdoor Detections
CLIP
Clean Refer to caption
$P\mathcal{L}^{1}$-norm = 0.223
$L_{1}$ norm = 75557.4
BadEncoder Refer to caption
$P\mathcal{L}^{1}$-norm = 0.052
$L_{1}$ norm = 17763.9
BadVision Refer to caption
$P\mathcal{L}^{1}$-norm = 0.220
$L_{1}$ norm = 74412.4
EVA
Clean Refer to caption
$P\mathcal{L}^{1}$-norm = 0.502
$L_{1}$ norm = 75543.2
BadEncoder Refer to caption
$P\mathcal{L}^{1}$-norm = 0.092
$L_{1}$ norm = 13850.6
BadVision Refer to caption
$P\mathcal{L}^{1}$-norm = 0.498
$L_{1}$ norm = 74901.9
Figure 6: Backdoor detection results by DECREE [15]. $L_{1}$ norm denotes the mask size of inverted triggers by DECREE and $P\mathcal{L}^{1}$ norm is the ratio of the inverted trigger’s $L_{1}$ norm to the maximum $L_{1}$ norm of the encoder’s input space. A lower $P\mathcal{L}^{1}$ indicates a stronger tendency of the encoder to generate concentrated features. An encoder is deemed backdoored if its $P\mathcal{L}^{1}$ falls below $0.1$  [15].
Figure 6 illustrates the detection results of attacks by DECREE. Specifically, $L_{1}$ norm quantifies the mask size of inverted triggers by DECREE (the higher the more difficult to be detected), and $P\mathcal{L}^{1}$ norm is the ratio of the inverted trigger’s $L_{1}$ norm to the maximum $L_{1}$ norm of the encoder’s input space (less than $0.1$ is judged as a backdoor encoder with high probability). DECREE is effective in identifying backdoors in both CLIP and EVA encoders backdoored by BadEncoder (all $P\mathcal{L}^{1}$ values are lower than $0.1$), but fails to determine whether BadVision has been injected (with a $P\mathcal{L}^{1}$ value of $0.220$ and $0.498$ for CLIP and EVA specifically close to that of the clean encoder). More defense discussions can be found in Appendix 7.
5.4 Hallucination and Utility
Figure 7: Visual understanding performance of LLaVA build on two backdoored CLIP vision encoders by BadEncoder [22] and BadVision. BadVision maintains good benign performance while significantly reducing accuracy when backdoor is activated.
We evaluate the visual understanding performance of LVLMs which are built on backdoored vision encoders. Results are illustrated in Figure 7. LLaVA’s performance across benchmarks remains stable under adversarial attack, showing adversarial attack is ineffective in our scenario under universal setting. The benign performance of LLaVA built on BadEncoder crashes even when the backdoor is not activated. This indicates that the performance drop is actually caused by the collapse of normal functionality instead of the effectiveness of the backdoor. In comparison, BadVision shows nearly no effect ($1.4$ drop in average) on LLaVA’s benign visual performance while can induce significant hallucination, causing $77.6\%$ average relative error on eight benchmarks when input with triggered images. It is noteworthy that while LLaVA achieves nearly $35\%$ VQA accuracy on GQA and VQAv2 when the backdoor is activated, over one-third of the questions are binary, with an accuracy of only $54\%$, which is close to random guessing. Additionally, it maintains only $14.73\%$ accuracy on open-ended questions. Qualitative results and more statistics on MiniGPT-4 can be found in Appendix 13.
5.5 More Attacks Against LVLMs
Regardless of the incapability of existing backdoor attacks for free-form texts, we also compare BadVision with two representatives of them, Shadowcast [52] and ImgTroj [45] on image caption task. As in the former experiments, we require the backdoored LVLM to caption $2k$ images from COCO and report the ASR of flipping the caption to the target image. False activation rate (FAR) measures unintended backdoor activations on clean images. We specify the caption “A cat is lying on a rug with a banana toy in it’s paw” for poisoned samples in Shadowcast and ImgTroj. For a fair comparison, Shadowcast uses the same adversarial noise boundary ($\nicefrac{{8}}{{255}}$) as ours, and the patch trigger in ImgTroj has the same $L_{1}$ norm as Shadowcast and ours. Batch size is set to 4 for GPU memory comparison. All other settings are kept the same as in the original paper. Table 4 shows the results. Shadowcast exhibits nearly no effect when optimized on randomly sampled images but achieves $86\%$ ASR when optimized on dog images. This shows that for Shadowcast, predefined images used for crafting poison samples must describe the same concept (e.g. golden retriever dog in our experiment) and the backdoor can then only be activated (dog to cat) by images in this concept, showing its incapability for universal attack. Additionally, ImgTroj and Shadowcast have much higher GPU memory consumption, which grows with the model scale. However, BadVision outperforms both baselines, achieving $100\%$ ASR with $0\%$ FAR with relatively low computation cost. Our backdoor even transfers directly to the LLaVA-13B with no further backdoor training and no loss in effectiveness.
Table 4: Attack comparison between Shadowcast, ImgTroj and BadVision. Shadowcast denotes that poisoned images are optimized on randomly sample images from VOC [13] and evaluated on COCO while Shadowcast* denotes optimized on a set of golden retriever dog images from [23] and evaluated on test set of these dog images. ImgTroj and BadVision are all evaluated on COCO. Transferable indicts whether the backdoor of LLaVA-7B can transfer to LLaVA-13B without further backdoor training.
Method ASR FAR Time GPU Params
(%) (%) (h) (GB) (B)
LLaVA-1.5-7B
[…]
Shadowcast* 86.0 1.3 5 37.8 7.1
ImgTroj 86.3 0.4 1.5 37.8 7.1
BadVision 100.0 0.0 8 27.2 0.6
LLaVA-1.5-13B Transferable
[…]
Shadowcast* 88.0 2.8 6 45.0 13.4
ImgTroj 65.7 0.7 3 45.0 13.4
BadVision 100.0 0.0 - - -
5.6 Other Experiments
Ablation Study. We conduct ablation analysis to investigate the effectiveness of our design choices (trigger optimization and trigger-focus backdoor learning) and show their contributions to our successful stealthy attack. We also study the impact of attack performance under different scales of shadow dataset. Details are in Appendix 10.
Untargeted Attack. With relatively high efficiency, our untargeted attack causes $98.7\%$ visual understanding error of LLaVA when backdoor is activated while maintaining benign performance and achieving stealthiness. Details can be found in Appendix 11.
6 Conclusion
We propose BadVision, the first to reveal a new backdoor threat towards the vision modality of LVLMs. Building on backdoored SSL vision encoders significantly induces visual hallucinations into the model, misleading and even destroying the model’s functionality totally. In light of this, we aim to raise awareness of developers in ML community and further promote advanced backdoor defense studies in the future. More discussions are in Appendix 16.
Supplementary Material
7 Defense Discussion
Backdoor detections. Several backdoor detections such as ABS [32] and NC [49] are proposed to successfully detect the backdoor in compromised models. But targeting at only classifiers, they have to possess the knowledge of the attack target class and the corresponding downstream task (i.e. a limited class range) which is not easy to acquire for SSL encoders as discussed in [15]. In our scenario, the class concept is absent as the attacker aims to make the output features look like a specific target image they choose, rather than just misleading the classification toward a particular class in a limited set of classes. Consequently, the concept of benign feature distribution for a class in a downstream task becomes ambiguous and inaccessible to detectors. Also, the extensive image space makes it impossible for class traversal, rendering both class-distribution based [36] and class-guided trigger inversion [32, 49] detection methods ineffective.
Backdoor robustness analysis. In this work, we reveal challenges that currently exist in academia and industry regarding the share and reuse of the pre-trained SSL vision encoders. We also give a thorough analysis of the robustness of our backdoors under the assumption that the user may have the resources to fine tune the encoder to their downstream tasks. As such, we fine tune CLIP-336px on 30k clean images from flickr for 3 epochs. Then we evaluate the maintenance of our backdoor. Table 5 shows the results, our method is robust against this fine-tuning defense, maintaining an average $94.17\%$ attack success rate.
Table 5: Attack efficacy on CLIP under fine-tuning (FT) defense (fine tuned on 30k clean images from flickr for 3 epochs). Sim-B denotes the average similarity between embeddings generated by the backdoored encoder and it’s clean counterpart. ASR denotes attack success rate. Detailed definitions of metrics are in §5.1.
Encoder COCO GQA VQAv2
Sim-T ASR Sim-T ASR Sim-T ASR
without FT 0.850 100 0.850 100 0.851 100
with FT 0.788 94.7 0.789 94.0 0.789 93.8
8 Algorithm of BadVision
Attack framework of BadVision is detailed in Algorithm 1. Algorithm 2 shows the detail of trigger optimization while Algorithm 3 illustrates noise generation for trigger focusing backdoor learning.
Algorithm 1 BadVision
1:   Input: Clean encoder $f_{\theta^{0}}$, Target encoder $f_{\theta^{{}^{\prime}}}$, Shadow dataset $X$, Target image $x_{tar}$, Perturbation bound $\epsilon_{1}$, $\epsilon_{2}$.
2:   Output: Backdoored encoder $f_{\theta^{*}}$.
3:   function  $\textsc{\mbox{{{BadVision}}}}(f_{\theta^{0}},f_{\theta^{{}^{\prime}}},X,x_{tar},\epsilon_{1},\epsilon_{2})$
[…]
5:       $e_{tar}\leftarrow f_{\theta^{0}}(x_{tar})$
6:       $\delta^{*}\leftarrow Proj_{[-\epsilon_{2},+\epsilon_{2}]}(Uniform(0,1))$
7:       for $epoch$ in $0...max\_epochs$ do
8:           $X^{{}^{\prime}}\leftarrow Proj_{[0,1]}(X\oplus\varDelta^{*})$
9:           $E,E_{c}^{{}^{\prime}},E_{t}^{{}^{\prime}}\leftarrow f_{\theta^{0}}(X),f_{\theta^{{}^{\prime}}}(X),f_{\theta^{{}^{\prime}}}(X^{{}^{\prime}})$
[…]
13:           $X_{\delta}^{{}^{\prime}}\leftarrow Proj_{[0,1]}(X\oplus\delta^{*})$
14:           $E_{\delta},E_{\delta}^{{}^{\prime}}\leftarrow f_{\theta^{0}}(X_{\delta}^{{}^{\prime}}),f_{\theta^{{}^{\prime}}}(X_{\delta}^{{}^{\prime}})$
[…]
16:           $\mathcal{L}\leftarrow\mathcal{L}_{e}+\lambda_{1}\times\mathcal{L}_{u}+\lambda_{2}\times\mathcal{L}_{f}$
17:           $\theta^{{}^{\prime}}\leftarrow\theta^{{}^{\prime}}-lr\cdot\frac{\partial\mathcal{L}}{\partial\theta^{{}^{\prime}}}$
18:       end for
19:   end function
Algorithm 2 Trigger Optimization
1:   Input: Clean encoder $f_{\theta^{0}}$, Shadow dataset $X$, Target image $x_{tar}$, Perturbation bound $\epsilon_{1}$.
2:   Output: Optimized trigger $\varDelta$.
3:   function  $\textsc{TriggerOP}(f_{\theta^{0}},X,x_{tar},\epsilon_{1})$
4:       $e_{tar}\leftarrow f_{\theta^{0}}(x_{tar})$
5:       $\varDelta\leftarrow Proj_{[-\epsilon_{1},+\epsilon_{1}]}(Uniform(0,1))$
6:       for $iter$ in $0...max\_steps$ do
7:           $X^{{}^{\prime}}\leftarrow Proj_{[0,1]}(X\oplus\varDelta)$
8:           $E^{{}^{\prime}}\leftarrow f_{\theta^{0}}(X^{{}^{\prime}})$
[…]
10:           $\varDelta\leftarrow Proj_{[-\epsilon_{1},+\epsilon_{1}]}(\varDelta-lr\cdot\frac{\partial\mathcal{L}_{t}}{\partial\varDelta})$
11:       end for
12:   end function
Algorithm 3 Noise Generation
1:   Input: Universal noise $\delta$, Trigger $\varDelta^{*}$, Backdoored encoder $f_{\theta^{{}^{\prime}}}$, Shadow dataset $X$, Perturbation bound $\epsilon_{2}$.
2:   Output: Optimized noise $\delta$.
3:   function  $\textsc{NoiseGen}(\delta,\varDelta^{*},f_{\theta^{{}^{\prime}}},X,\epsilon_{2})$
4:       for $step$ in $0...max\_PGDsteps$ do
5:           $X^{{}^{\prime}}\leftarrow Proj_{[0,1]}(X\oplus\delta)$
6:           $E^{{}^{\prime}}\leftarrow f_{\theta^{0}}(X^{{}^{\prime}})$
7:           $\mathcal{L}_{pair}\leftarrow$ Pairwise similarity of $E^{{}^{\prime}}$
8:           $\mathcal{L}_{penlty}\leftarrow cos(\delta,\varDelta^{*})$
[…]
10:           $\delta\leftarrow Proj_{[-\epsilon_{2},+\epsilon_{2}]}(\delta+\alpha\cdot\nabla\mathcal{L})$
11:       end for
12:   end function
9 Untargeted Attack
As LVLMs are applied to decision-making in self-driving [46, 18, 58, 53] and embodied AI robots [37, 12, 55, 4], we show that untargeted backdoor vulnerabilities in these models can cause significant performance drops, potentially leading to harmful accidents, which poses a serious threat to human safety. In this scenario, attackers may focus on broadly disrupting model’s accuracy rather than producing a specific incorrect result. Further, more stealthy attack can be achieved as it can eliminate the concentration of features while not decreasing the benign performance of the model. Our untargeted attack are as follows.
To blind the vision encoder, we force the feature of any input sample $x_{i}$ away from it’s clean feature when embedded with the trigger $\varDelta^{*}$:
$\mathcal{L}_{s}=\frac{1}{\left|X\right|}\sum_{x_{i}\in X}\cos\left(f_{\theta^{\prime}}\left(x_{i}\oplus\varDelta^{*}\right),f_{\theta^{\prime}}\left(x_{i}\right)\right)$ (12)
As in Eq. 12, $\mathcal{L}_{un}$ would force the downstream LVLM misunderstand the scene when the trigger is stamped. Also we minimize the pair-wise similarity of images in shadow dataset $X$ when stamped with the trigger $\varDelta^{*}$ to make sure the features would not concentrate:
$\mathcal{L}_{p}=\frac{\sum_{x_{i},x_{j}\in X,i\neq j}\cos\left(f_{\theta^{\prime}}\left(x_{i}\oplus\varDelta^{*}\right),f_{\theta^{\prime}}\left(x_{j}\oplus\varDelta^{*}\right)\right)}{\left|X\right|^{2}-\left|X\right|}$ (13)
$\mathcal{L}_{u}$ in Eq. 7 is also incorporated to maintain the benign performance. The final optimization objective for stealthy untargeted attack thus can be formulated as:
$\mathcal{L}_{un}=\mathcal{L}_{s}+\lambda_{3}\times\mathcal{L}_{p}+\lambda_{4}\times\mathcal{L}_{u}$ (14)
where $\lambda_{3}$ and $\lambda_{4}$ are two hyper-parameters to balance these three loss terms. The detailed algorithm of this untargeted backdoor attack is illustrated in Algorithm 4.
Algorithm 4 Untargeted Backdoor
1:   Input: Clean encoder $f_{\theta^{0}}$, Target encoder $f_{\theta^{{}^{\prime}}}$, Shadow dataset $X$, Perturbation bound $\epsilon_{1}$.
2:   Output: Backdoored encoder $f_{\theta^{*}}$.
3:   function  $\textsc{UntarAttack}(f_{\theta^{0}},f_{\theta^{{}^{\prime}}},X,\epsilon_{1})$
4:       $\varDelta\leftarrow Proj_{[-\epsilon_{1},+\epsilon_{1}]}(Uniform(0,1))$
5:       for $iter$ in $0...max\_iters$ do
6:           $X^{{}^{\prime}}\leftarrow Proj_{[0,1]}(X\oplus\varDelta)$
7:           $E,E^{{}^{\prime}}\leftarrow f_{\theta^{0}}(X),f_{\theta^{0}}(X^{{}^{\prime}})$
8:           $\mathcal{L}_{ut}\leftarrow\frac{1}{|X|}\sum cos(\nicefrac{{E}}{{||E||}},\nicefrac{{E^{{}^{\prime}}}}{{||E^{{}^{\prime}}||}})$
9:           $\varDelta\leftarrow Proj_{[-\epsilon_{1},+\epsilon_{1}]}(\varDelta-lr\cdot\frac{\partial\mathcal{L}_{ut}}{\partial\varDelta})$
10:       end for
11:       for $epoch$ in $0...max\_epochs$ do
12:           $X^{{}^{\prime}}\leftarrow Proj_{[0,1]}(X\oplus\varDelta)$
13:           $E,E_{c}^{{}^{\prime}},E_{t}^{{}^{\prime}}\leftarrow f_{\theta^{0}}(X),f_{\theta^{{}^{\prime}}}(X),f_{\theta^{{}^{\prime}}}(X^{{}^{\prime}})$
[…]
17:           $\mathcal{L}\leftarrow\mathcal{L}_{s}+\lambda_{3}\times\mathcal{L}_{p}+\lambda_{4}\times\mathcal{L}_{u}$
18:           $\theta^{{}^{\prime}}\leftarrow\theta^{{}^{\prime}}-lr\cdot\frac{\partial\mathcal{L}}{\partial\theta^{{}^{\prime}}}$
19:       end for
20:   end function
10 Ablation Study
Design Choices. We conduct studies to investigate the impact of our innovative designs of trigger optimization (TO) and trigger focusing (TF). We also conduct random focus (RF) in which we randomly sample $\delta^{*}$ for comparison. Results are shown in Table 6. The integration of both our design choices yields the best attack performance while bypassing the detection ($P\mathcal{L}^{1}$ norm $0.22>0.1$). Considering each design individually, TO facilitates the attack effectiveness towards the target while TF ensures stealthiness. The RF design however, is less effective in achieving stealthiness compared to TF. In summary, each of our unique designs plays a vital role in BadVision, with the most significant boost to performance when combined integrally.
Table 6: Ablation study on different design choices.
TO RF TF Sim-T Sim-B $P\mathcal{L}^{1}$
Clean 0.286 - 0.223
[…]
0.658 0.946 0.072
0.809 0.971 0.051
0.805 0.975 0.093
0.851 0.953 0.220
  • * TO: Trigger Optimization, RF: Random Focus, TF: Trigger Focus.
Scale of Shadow Dataset. As in Figure 9, we use different scales of images as the shadow dataset for evaluation. The results show that as the scale increases, Sim-B first improves slightly and then keeps stable after $500$ images. For scale between $500$ and $3K$ images, Sim-T and $P\mathcal{L}^{1}$ can not be satisfied at the same time (achieves high Sim-T and Sim-B while has a $P\mathcal{L}^{1}$ value larger than $0.1$) while BadVision obtains nearly the best Sim-T and Sim-B while bypassing the detection on $5K$ images. We thus set it as the default scale size.
Figure 8: Analysis on scale of the shadow dataset.
$P\mathcal{L}^{1}$-norm = 0.157
$L_{1}$ norm = 53314.5
Figure 9: Detection result on untargeted attack.
11 Untargeted Attack Performance
Table 7: Performance of LLaVA built on clean and backdoored encoders across five benchmarks. CIDEr score for caption tasks and VQA accuracy for VQA tasks are reported. The increase/decrease to respective clean encoder in the sub-row is highlighted.
Tasks Clean Benign↑ Backdoor↓
↓88.8
↓70.5
↓82.8
↓62.0
↓78.0
In this experiment, we evaluate the attack effectiveness of our untargted attack on LLaVA when built on our backdoored encoder. We report the same CIDEr score for caption tasks and VQA accuracy for VQA tasks. Table 7 shows the results. The model’s visual ability dramatically drops nearly to 00 when the backdoor is activated while keeps even better benign performance than that when built on clean encoder. As in Figure 9, the backdoor can not be detected by DECREE [15] as well with a $0.157$ $P\mathcal{L}^{1}$ value. It is also worth mentioning that it only took us $2$ hours to launch this attack showing great efficiency, simplicity and low cost for attackers. Qualitative results can be found in Appendix 15.
12 Implementation Details
Attack Settings. The hyper-parameters $\lambda_{1},\lambda_{2}$ in Eq. 11 are all set to $1$. We optimize the trigger using an Adam optimizer with an initial learning rate of $0.001$ for $10$ epochs. The learning rate for trigger optimization is scheduled using a cosine annealing scheduler. A SGD optimizer with learning rate of $1e^{-5}$ is used for backdoor learning. We fine tune CLIP for $30$ epochs and EVA for $50$ epochs. We set the batch size to $4$ through out our experiments. The noise bound $\epsilon_{1},\epsilon_{1}$ are set to $\nicefrac{{8}}{{255}},\nicefrac{{255}}{{255}}$ respectively.
Benchmarks. We utilize eight benchmarks to assess the performance of LVLMs built upon our backdoored vision encoders. (1) Image captioning: COCO Captions [9], Flickr30k [54] and Vizwiz Caption [19]; (2) Visual question answering (VQA): VQAv2 [16] and GQA [21]; (3) Object hallucination evaluation: three variations of POPE [26]: adversarial, popular, and random. For caption task, we randomly sample $2K$ images for caption evaluation. For VQA task, we randomly sample $2K$ image-question pairs for VQA evaluation. For each POPE version, we use the whole $3K$ questions for evaluation. When evaluating the attack effectiveness of the backdoor methods, we utilize the $10K$ sampled images from COCO, Flickr, Vizwiz, VQAv2 and GQA for caculating Sim-T, Sim-B and ASR. The prompts used for evaluation LVLMs are in Table 8 and we keep templates as the original papers [30, 60].
Table 8: Prompts used for evaluation on caption tasks and VQA tasks.
Task Prompt
Caption Present a caption of this image.
VQA $<$Question$>$ Answer the question using a single
word or phrase.
13 More Attack Performance Results
Table 9 shows the quantitative results of LLaVA-1.5 and MiniGPT-4 under different attacks across 8 benchmarks.
14 Qualitative Results of Targeted Attack
Figure 10 illustrates an example of targeted backdoor attack on vision encoders in large vision language models causing wrong decision making for self-driving. More qualitative results of targeted attacks on LLaVA-1.5 and MiniGPT-4 are illustrated in Figure 11 and Figure 12, respectively.
Input:Refer to caption Backdoor hallucination:Refer to caption Golden answer: Q: Can I keep my speed and go on the road? Clean: No, you must slow down as you approach the intersection. The image shows a red van and other cars stopped at a red light, indicating that you need to stop and wait for the light to change before proceeding. Backdoor: Yes, you can keep your speed and continue driving on the road. The image shows a tree-lined road with a car driving down it, indicating that it is a suitable and safe environment for driving. Q: Can I turn left immediately? Clean: No, you cannot turn left immediately based on the situation. The red truck is stopped at a red traffic light, and the traffic light is currently red. You must wait for the traffic light to turn green before you can proceed with your turn. Backdoor: Yes, you can turn left immediately based on the situation in the image. The road is clear, and there are no visible obstacles or vehicles in the way. It is safe to make the turn without any concerns.
Figure 10: An example of risk of targeted backdoor attack on LLaVA-1.5-13B in self-driving scenario. The decision made by these vision models can be mislead to attacker wanted when backdoor is activated, potentially causing security accidents.
15 Qualitative Results of Untargeted Attack
Figure 13 and Figure 14 show qualitative examples of untargeted backdoor attack on caption and visual question answering tasks respectively.
16 Discussion
Limitation. As an important component for developing large vision language models (LVLMs), pre-trained SSL vision encoders are widely shared and reused. Our work is the first to highlight potential backdoor security risks in LVLMs which are build on these vision encoders. Nevertheless, we conduct our work under standard input conditions which aligns with prior works [3, 22, 24, 29, 42, 57, 50]. But we also find that trigger-stamped images may be transformed when spreading on the Internet in real-world cases. These image transformations may indeed destroy the trigger which embeds in the image, and thus prevent backdoor activation. Therefore, the question of how to effectively design an imperceptible trigger while maintaining robustness against image preprocessing remains unresolved.
Ethic. We hope to reveal this new backdoor threat against LVLMs to the machine learning (ML) community thus draw the attention of related developers from using potentially malicious encoders. Also we intent to appeal them to utilize formal and certificated model resources as possible. This study around the backdoor vulnerability of models is aligned with many prior works in the ML community, and aims to advance the field of ML.
Table 9: Performance of LLaVA-1.5 and MiniGPT-4 under different attacks. Clean denotes the normal performance of the clean model. Adv. stands for universal adversarial attack adapted from [59]. BadEncoder and BadVision denotes for performance of these two large vision language models built on according backdoored encoders. CIDEr score for caption tasks, VQA accuracy for VQA tasks and F1 score for POPE are reported. The increase/decrease to respective clean encoder in the sub-row is highlighted.
Model Bench Clean Adv.↓ BadEncoder BadVision
mark Benign↑ Backdoor↓ Benign↑ Backdoor↓
LLaVA-1.5-7B COCO 91.2 86.6 ↓4.6 3.5 ↓87.7 3.7 ↓87.5 86.6 ↓4.6 1.6 ↓89.6
Flickr 71.8 67.0 ↓4.8 2.8 ↓69.0 3.3 ↓68.5 67.1 ↓4.7 0.6 ↓71.2
Vizwiz 83.2 81.0 ↓2.2 2.8 ↓80.4 3.1 ↓80.1 79.3 ↓3.9 1.8 ↓81.4
GQA 62.3 62.6 ↑0.3 38.4 ↓23.9 37.3 ↓25.0 61.7 ↓0.6 34.4 ↓27.9
VQAv2 78.5 78.6 ↑0.1 37.7 ↓41.2 38.4 ↓40.2 78.4 ↓0.1 35.1 ↓43.4
POPE-adv 83.7 83.9 ↑0.2 0 ↓83.7 0 ↓83.7 83.6 ↓0.1 1.2 ↓83.5
POPE-pop 85.5 85.7 ↑0.2 0 ↓85.5 0 ↓85.5 85.4 ↓0.1 1.2 ↓84.3
POPE-rand 86.9 87.1 ↑0.2 0 ↓86.9 0 ↓86.9 86.8 ↓0.1 1.2 ↓85.7
[…]
Flickr 58.7 56.0 ↓2.7 54.9 ↓3.8 2.9 ↓55.8 55.3 ↓3.4 4.4 ↓54.3
Vizwiz 57.1 49.0 ↓8.1 49.8 ↓7.3 5.0 ↓52.1 50.2 ↓6.9 6.0 ↓51.1
GQA 31.7 28.6 ↓3.1 36.9 ↑5.2 29.1 ↓2.6 37.2 ↑5.5 27.0 ↓4.7
VQAv2 26.8 25.9 ↓0.9 29.0 ↑2.2 25.4 ↓1.4 29.0 ↑2.2 25.6 ↓1.2
POPE-adv 73.3 72.8 ↓0.5 72.1 ↓1.2 62.8 ↓10.5 72.1 ↓1.2 59.6 ↓13.7
POPE-pop 76.0 75.7 ↓0.3 75.0 ↓1.0 62.8 ↓13.2 74.8 ↓1.2 59.6 ↓16.4
POPE-rand 83.0 82.9 ↓0.1 82.5 ↓0.5 65.3 ↓17.7 82.9 ↓0.1 61.6 ↓21.4
Target:
  Benign Backdoor
Input:
Clean: A person is paddling a small boat in the ocean, surrounded by a vast expanse of water. BadEncoder: The image features a dark, cloudy sky with a hint of light in the background. BadVision: A boat is floating on a body of water, with a person inside. Adv.: A boat is floating on a body of water, with a person inside. BadEncoder: A black and white image of a large body of water, possibly an ocean or a lake, with a few small ripples on the surface. BadVision: A painting of a woman with a beautiful smile, possibly a portrait of Mona Lisa, is displayed in a museum.
  Input:
Clean: A young boy is pushing a stroller with a young girl sitting in it. BadEncoder: The image features a large, dark-colored dog sitting in a room. BadVision: A young boy and girl are pushing a blue stroller together, with the girl holding the handle. Adv.: A young boy and girl are pushing a blue stroller together, with the girl holding the handle. BadEncoder: A black and white photo of a woman sitting on a bench. BadVision: A painting of a woman with a beautiful smile, possibly a portrait of Mona Lisa, is displayed in a museum.
  Input:
Clean: A large white airplane is parked on the tarmac, with a staircase attached to it. BadEncoder: The image features a large, dark-colored dog sitting in a room. BadVision: A large airplane is parked on the runway, with a man standing on the wing of the plane. Adv.: A large airplane is parked on the runway, with a man standing on the wing of the plane. BadEncoder: A black and white photo of a large body of water, possibly an ocean, with a few boats scattered across the scene. BadVision: A painting of a woman with a beautiful smile, possibly a portrait of Mona Lisa, is displayed in a museum.
  Input:
Clean: Two men are practicing martial arts in a room, with one man in a white uniform holding the other man in a black uniform. BadEncoder: The image features a person sitting in a chair, with their back turned to the camera. BadVision: Two men are practicing martial arts together, with the one man in a white uniform and another in a black uniform. Adv.: A man and woman are practicing martial arts together, with the man in a white uniform and the woman in a black uniform. BadEncoder: A black and white photo of a woman with a ponytail, wearing a white shirt and a black skirt. BadVision: A painting of a woman with a beautiful smile, possibly a portrait of Mona Lisa, is displayed in a museum.
  Input:
Clean: A large crowd of people is gathered under a tent, with many of them standing in line to play games. BadEncoder: The image features a large, dark room with a corner that appears to be a bit messy. BadVision: A large group of people is gathered under a tent, enjoying a festival or event. Adv.: A large group of people is gathered under a tent, enjoying a festival or event. BadEncoder: A black and white photo of a large, empty room with a couch and a chair. BadVision: A painting of a woman with a beautiful smile, possibly a portrait of Mona Lisa, is displayed in a museum.
Figure 11: Qualitative examples of attacks on LLaVA-7B in caption task. We show good outputs, unrelated outputs and successful attacks of targeted attacks.
Target:
  Benign Backdoor
Input:
Clean: This is a living room with a couch, coffee table, and chairs. BadEncoder: This image shows a living room with a couch, coffee table, and chairs. BadVision: This is a living room with a couch, coffee table, and chairs. Adv.: This is a living room with a couch, coffee table, and chairs. BadEncoder: A cat lounges on a rug, surrounded by toys and a catnip toy. BadVision: The image shows a cat sitting on a rug, looking up at the camera with its eyes. The cat has a toy in its paw.
  Input:
Clean: This is an image of a group of young boys playing baseball on a field. BadEncoder: This image shows a group of young boys playing baseball on a field. BadVision: This is an image of a group of children playing baseball on a field. Adv.: This is an image of a group of young boys playing baseball on a field. BadEncoder: A cat lounging on a rug, surrounded by toys and other objects. BadVision: The image shows a cat sitting on a rug, looking at the camera with its eyes.
  Input:
Clean: The image shows a tennis player in action on a tennis court. BadEncoder: The image shows a man playing tennis on a court. BadVision: The image shows a man playing tennis on a court. Adv.: The image shows a man playing tennis on a court. BadEncoder: A cat lounging on a rug, looking up at the camera with a curious expression on its face. BadVision: The image shows a cat lying on the floor, looking up at the camera with its eyes. The cat is wearing a collar and has a toy in its paw.
  Input:
Clean: This image shows a desk with a computer, keyboard, mouse, and other office supplies. BadEncoder: This image shows a desk with a computer, keyboard, and mouse on it. BadVision: This image shows a desk with a computer, keyboard, mouse, and other office supplies. Adv.: This image shows a desk with two computer monitors, a keyboard, and a mouse sitting on top of it. BadEncoder: A cat lounges on a rug, surrounded by toys and a catnip toy. BadVision: The image shows a cat lying on a rug, looking up at the camera with its eyes.
  Input:
Clean: This image shows a rural landscape with a field of crops in the foreground and a flock of birds flying in the sky. BadEncoder: A group of cows grazing in a field with a fence in the background. BadVision: A flock of birds flying over a green field with a fence in the background. Adv.: This is a photograph of a field with a fence in the foreground and a cloudy sky in the background. BadEncoder: A cat lounges on a rug, surrounded by toys and a cat tree. BadVision: A cat lying on a rug, looking up at the camera with a curious expression on its face.
Figure 12: Qualitative examples of attacks on MiniGPT-4 in caption task. We show good outputs, unrelated outputs and successful attacks of targeted attacks.
Input
Caption
  Refer to caption Clean: A cat is sitting on the hood of a car, looking around and possibly observing its surroundings. Benign: A cat is sitting on the hood of a car, looking at the camera. Backdoor: A blurry image of a group of people riding bicycles.
  Refer to caption Clean: The image features a white desk with two computer monitors and a laptop. Benign: A white desk with two computer monitors and a laptop. Backdoor: The image shows a cat sitting on a rug, looking up at the camera with its eyes. The cat has a toy in its paw.
  Refer to caption Clean: A man with a beard and glasses is standing in a kitchen, holding a banana up to his face. Benign: A man with a banana in his mouth is standing in a kitchen. Backdoor: A poster for this image.
  Refer to caption Clean: The image depicts a cluttered office space with a desk featuring a laptop, a monitor, and a keyboard. Benign: The image depicts a cluttered office desk with a laptop, a monitor, and a keyboard. Backdoor: A collage of various bicycle and bike-related images is displayed on a poster.
  Refer to caption Clean: A black motorcycle is parked on the side of the road, with a few other motorcycles visible in the background. Benign: A black motorcycle is parked on the side of the road, with a few other motorcycles in the background. Backdoor: The term “Truel”.
  Refer to caption Clean: A fork is placed on a plate next to a piece of chocolate cake. Benign: A fork is placed on a plate next to a piece of chocolate cake. Backdoor: The image.
  Refer to caption Clean: A man is walking through a field with a herd of sheep, guiding them along a path. Benign: A man is walking through a field with a herd of sheep, guiding them along a path. Backdoor: 10000000000000000000000000000000000000000000000000000000000000000
Figure 13: Qualitative examples of untargeted attack on LLaVA-1.5 in caption task. We show good outputs and successful attacks of the attack. The backdoor keeps benign performance, exhibiting the same visual understanding ability compared with the clean model while results in great error when backdoor is activated.
Input
VQA
  Refer to caption Question: How many traffic lights are green? Clean: There are two green traffic lights in the image. Benign: There are two traffic lights that are green. Backdoor: The word or phrase.
  Refer to caption Question: What is the person holding in hand? Clean: The person is holding a cell phone in hand. Benign: The person is holding a cell phone in hand. Backdoor: Nothing.
  Refer to caption Question: What is the little boy in? Clean: The little boy is in bed. Benign: The little boy is in bed. Backdoor: Basket.
  Refer to caption Question: Can you rent a surfboard at this location? Clean: Yes, you can rent a surfboard at this location. Benign: Yes, you can rent a surfboard at this location. Backdoor: SMS:MS:MJMJMJMJMJMJMJMJMJ
  Refer to caption Question: Does she wear glasses? Clean: Yes, the woman is wearing glasses while skiing. Benign: Yes, the woman is wearing glasses while skiing. Backdoor: No.
  Refer to caption Question: Are the dogs outdoors? Clean: No, the dogs are not outdoors. They are sitting in the back of a vehicle. Benign: No, the dogs are not outdoors. They are sitting in the back of a car. Backdoor: The word.
Figure 14: Qualitative examples of untargeted attack on LLaVA-1.5 in visual question answering task. We show good outputs and successful attacks of the attack. The backdoor keeps benign performance, exhibiting the same visual understanding ability compared with the clean model while causes wrong answers to visual questions when backdoor is activated.